OpenStudy (anonymous):

Last question of the night! find the limit as h approaches 0: [cos(x+h)^7 -cosx^7]/h

OpenStudy (anonymous):

This may seem simple, but I just can't figure out where to start. Brain, don't quit on me now!

OpenStudy (anonymous):

It's multiple choice: a. -7x^6sinx^7 b. -7cos^6xsinx c. -sin7x^13 d. infinity e. 0

OpenStudy (anonymous):

@phi are you still on? could you please help me once more?

OpenStudy (anonymous):

I know I can't use direct substitution because that produces indeterminate form. yes. but I don't think they want me to factor or simplify, either. am I supposed to do something with differentiation here?

OpenStudy (phi):

can you use L'Hospital's Rule ?

OpenStudy (anonymous):

I've heard of that! but strangely, it doesn't come up as a lesson until a few units later, at least. I wasn't sure if that is how they wanted me to solve it then. I have never worked with this rule. could you please help with that?

OpenStudy (anonymous):

it's like, differentiate numerator and then differentiate denominator and then take the limit, right?

OpenStudy (phi):

yes. you use it when you have an indeterminate form such as 0/0 (the case here) or infinity/infinity

OpenStudy (anonymous):

ok so how should I go about it when there are variables x and h? (oh dear, is that a question I should know the answer to?) >_<

OpenStudy (turingtest):

can we just identify this as the definition of the derivative of cos^7(x), or do we have to do this the long painful way?

OpenStudy (phi):

in this case, we treat x as a constant

OpenStudy (phi):

btw, just to make sure this is \[ \cos\left( (x+h)^7\right) \]?

OpenStudy (anonymous):

well, I can't say for sure, but i'm leaning towards long and painful. yes phi, it is

OpenStudy (anonymous):

so then we are differentiating with respect to h? is that what you mean by treating x as a constant?

OpenStudy (phi):

maybe it is better to keep track of x and get factors of dx/dh they will subtract out...

OpenStudy (turingtest):

if you look at the form, it's the definition of the derivative of cos(x^7), so you could just observe that and take the derivative wrt x. I really wouldn't know how to approach the problem otherwise, except to start hacking away with various trig identities or l'hospital. If anyone has a more rigorous proof I'd sure like to see it though :D

OpenStudy (anonymous):

I think we were going to use l'hospital. i'm still a bit confused though.

OpenStudy (phi):

turingtest's idea makes sense. We could use l'hospital, but if you have not studied it yet, maybe they want you to say: that is the definition of the derivative of cos(x^7). and find the derivative of cos(x^7)

OpenStudy (anonymous):

ok that does make sense but for the second part, are you saying that the whole equation is just an expanded definition of the derivative of cos(x^7)? so we would just take the derivative of that? (would we use chain rule?) u= cosx y= u^7 ?

OpenStudy (phi):

cos(u) u = x^7

OpenStudy (anonymous):

so derivative is: -sin(u)? -sinx^7?

OpenStudy (anonymous):

or would you take derivative of u before substituting back in? 7x^6 -7sinx^6?

OpenStudy (anonymous):

@phi are you there? which of those two would it be?

OpenStudy (phi):

d cos(u)= -sin(u) du u is x^7 du = d x^7 = 7 x^6 we get - sin u du = -sin(x^7) 7 x^6 or -7 x^6 sin(x^7)

OpenStudy (anonymous):

oh, okay! I missed a part. so then our answer is a! and no need for the l'hospital rule!

OpenStudy (phi):

no need for l'hospital btw, when I say d cos(u) that is short of d/dx cos (u) = -sin(u) du/dx when we get to d/dx x^7 we get 7 x^6 dx/dx or just 7 x^6

OpenStudy (anonymous):

right, ok! I believe that's it, then! I'll go over this once more to make sure I get it. thanks so much for your help today! i'm closing the question.